The sum of three numbers is 165. When the smallest number is multiplied by 7, the result is $n$. The value $n$ is obtained by subtracting 9 from the largest number. This number $n$ also results by adding 9 to the third number. What is the product of the three numbers?

Answer by ikleyn(52788)

.
The sum of three numbers is 165. When the smallest number is multiplied by 7, the result is $n$. The value $n$ is obtained by subtracting 9 from the largest number. This number $n$ also results by adding 9 to the third number. What is the product of the three numbers?
------------------------------------------------------------------

You use terms "smallest number", "largest number" and "third number".
They are not the terms of one semantic row. Which is bad.

Nevertheless, I conclude the "third number" is the "medium/middle number".

If it is not so, then your formulation is simply provocation.

From the condition, this system of three equations for three unknowns arises

s + m + l = 165,   (1)
7s = l - 9,        (2)
l - 9 = m + 9.     (3).

where s is the smallest, l is the largest and m is the medium ("the third") number. 

Solve it, if you can. 

If not, then confirm that the "third number" is the medium.

Or re-edit and re-post your quest.